20dm 3 so 2 diffuse through an aperture in 60sec. What vol O 2 will diffuse under similar condition in 30 sec?

20dmso diffuse through an aperture in 60sec. What vol O2 will diffuse under similar condition in 30 sec? 

Grade:12th pass

1 Answers

Vikas TU
4 years ago
 Accordding to Graham’s law of diffusion -
r1/r2 = sqrt(M1/M2)
r = Rate of diffusion of gas  M is molr mass 
20/60 * 30/v = sqrt(32/64)
v = 14.14 l 
Hope this helps